OpenStudy (anonymous):

Simplify the expression. Cubed root of 27(x-2)^3.

OpenStudy (anonymous):

\[\sqrt[3]{27(x-2)^3}= \sqrt[3]{3^3*(x-2)^3}\] in general cube root(y^3) = y so \[\sqrt[3]{3^3(x-2)^3}=3*(x-2)\]

OpenStudy (anonymous):

So it would not be 27(x-2) because you can simplify the 27, correct?

OpenStudy (anonymous):

yes because you use 3^3

OpenStudy (anonymous):

Does the cubed number and the power number cancel each other out?

OpenStudy (anonymous):

yes, cube root means a power of 1/3. so when you have cube root of 3^3, you have 3^(3/3) = 3^1 = 3

OpenStudy (anonymous):

Sorry I suck at math....

OpenStudy (ivancsc1996):

The answer is actually:\[3x-6\]:)

OpenStudy (anonymous):

How do you get 3x-6?

OpenStudy (ivancsc1996):

\[3(x-2)=3(x)+3(-2)=3x-6\] I used distribution which is basically\[a(b+c)=ab+ac\]

OpenStudy (ivancsc1996):

Get it or still having trouble with it?

OpenStudy (anonymous):

Not using the distributive property. It's the cubed root to the third power. Unless I'm throughly confused, which is probably the case....

OpenStudy (ivancsc1996):

Ok, the thing is that I am using the distributive to the answer 3(x-2) not the whole thing. This is the whola thing:\[\sqrt[3]{27(x-2)^{3}}=\sqrt[3]{27}\sqrt[3]{(x-2)^{3}}=3(x-2)=3(x)-3(2)=3x-6\]Remember that a cubed root is distributive can be broken down in multiplications and that a cubed power cancels a squared root.

OpenStudy (anonymous):

ok...thank you ....finally got it

OpenStudy (ivancsc1996):

No prob